OpenStudy (gibbs):

1. 1/10 + 1/4= 2. 1/6 + 1/4= 3. 1/6 + 5/8= 4. 4/9 + 4/15= 5. 5/6 - 2/9= 6. 2/5 +1/4= 7. 3/8 + 3/10= 8. 3/4 + 1/3= 9. 2/5 - 1/4= 10. 7/9 - 3/4= 11. 7/10 - 1/4= 12. 9/10 - 5/8= 13. 1/9 + 1/7= 14. 11/12 - 3/8= 15. 5/8 -1/6=

OpenStudy (gibbs):

Show worck plz

OpenStudy (gibbs):

medal!

OpenStudy (cwrw238):

to do all these you need to find the LCM of the 2 denominators before you can add or subtract LCM = Lowest Common Multiple ie the lowest number into which each of the 2 denominators will divide into exactly

OpenStudy (cwrw238):

for example find LCM of 4 and 6 first multiply:- 4*6 = 24 now 24 is a multiple of 4 and 6 but may not be the lowest multiple try dividing by 2 24 / 2 = 12 - this is your LCM - both 4 and 6 will divide into it and its the lowest multiple

OpenStudy (cwrw238):

so 1/4 + 1/6 - first convert these fractions to something / 12 to do this we multiply the 1 and 4 of first frsction by 3:- = 3/12 now multiply 1 and 6 of the 1/6 by 2: = 2/12 so we have 2/12 + 3/12 = 5/12 we can add the 2 and 3 here because the denominator 12 is same for both fractions

OpenStudy (gibbs):

ok next

OpenStudy (cwrw238):

so the anwser to no 2 is 5/12 follow the above process and you can do all 15

OpenStudy (cwrw238):

what do think the LCM of 4 and 10 is ( QUestion No. 1)

OpenStudy (gibbs):

1 is completed i need 7 through 15

OpenStudy (cwrw238):

4 * 10 = 40 is it 40 or 40/2 = 20?

OpenStudy (cwrw238):

ok No 7 what is the LCM of 8 and 10?

OpenStudy (cwrw238):

8 * 10 = 80 - now divide by 2 to see if there is a lower multiple into which 8 and 10 will divide

OpenStudy (gibbs):

80

OpenStudy (gibbs):

6/80

OpenStudy (cwrw238):

no - divide 80 by 2

OpenStudy (gibbs):

40

OpenStudy (cwrw238):

= 40 right now will 8 and 10 divide into 40?

OpenStudy (gibbs):

3/40

OpenStudy (cwrw238):

wher did you get the 3/40 from ? all im asking is do you think 40 is the LCM of 8 and 10?

OpenStudy (gibbs):

its the answer

OpenStudy (cwrw238):

3/10? is the answer for no 7?

OpenStudy (gibbs):

no

OpenStudy (cwrw238):

3/40?

OpenStudy (gibbs):

yes

OpenStudy (cwrw238):

no - no where near

OpenStudy (gibbs):

all i need now is 13 14 15

OpenStudy (texaschic101):

even if you do not get the LOWEST common denominator, you can still use it and then reduce at the end

OpenStudy (gibbs):

yes

OpenStudy (cwrw238):

40 is the LCM so you need to convert 3/8 to something over/40 this way: